OpenStudy (anonymous):

Given the function f(x) = 2(x + 10), find x if f(x) = 24.

OpenStudy (anonymous):

Put f(x) = 24 \[2(x + 10) = 24\] Divide both the sides by 2 and then find x..

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(x) = 2(x + 10) = 24}\) \( \color{Black}{\Rightarrow 2(x + 10) = 24 }\) \( \color{Black}{\Rightarrow x + 10 = 12 }\)

OpenStudy (anonymous):

17

OpenStudy (anonymous):

No check your answer again..

OpenStudy (anonymous):

7

OpenStudy (anonymous):

\[x + 10 = 12\] Subtract 10 from both the sides, and let us see what you get..

OpenStudy (anonymous):

68

OpenStudy (anonymous):

Oh My God.. When you subtract 12 - 10 then how much you get??

hero (hero):

f(x) = f(x) 2(x+10) = 24 2x + 20 = 24 2x = 24 - 20 2x = 4 x = 4/2 x = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Yes now you are correct..

Parth (parthkohli):

Yeah